# Instabilities of Gauged Q-Balls

### v1

@@ I Introduction
The stability of these solitons has been discussed in many papers, and the picture is still unclear. For global Q-balls, it has been shown Lee:1991ax ; Friedberg:1976me ; Paccetti:2001uh that the Q-balls are stable as long as ${d\omega\over dQ}<0$ (these quantities are defined below). For gauged Q-balls, the analogous statement is still believed to be true, but this has not been proven Panin:2016ooo .
Recently, a numerical analysis of the stability of gauged Q-balls was carried out Kinach:2022jdx . It was found that gauged Q-balls were unstable for certain parameters, and in particular were unstable to modes which had an angular dependence; indeed, for one particular parameter choice that was studied (denoted P2 in that paper), the instability had $L=4$. On the other hand, for other choices of parameters, the corresponding Q-balls were stable. In general, Q-balls with ${d\omega\over dQ}<0$ were stable.
@@
We then show in section V that the instability that we find has many characteristics in common with the numerical analysis of Kinach:2022jdx . In particular, we indeed find that for the parameter choice P2 in Kinach:2022jdx , the $L=4$ mode is unstable. Similarly, we find that for another parameter choice P1, the instability is absent, in agreement with the numerical analysis. We also find that Q-balls with ${d\omega\over dQ}<0$ are stable.
@@ II Q-Balls: Overview and Notation
We review here the basic properties of Q-balls; we will mostly follow the notation of Heeck:2020bau ; Heeck:2021zvk .
@@ III Spherical gauged Q-balls
This section follows the analysis of Heeck:2021zvk .
Large spherical Q-balls are described by a scalar profile that is approximately a step function;
$f^{(0)}\to\left\{\begin{array}[]{c}f_{+}\qquad\rho<R\\ 0\qquad\rho>R\end{array}\right.$ (12)
More precisely, we can divide the Q-ball into three regions.
In the interior $\rho\ll R$, the scalar is extremely close to a solution of the RHS of (5) i.e. $f^{(0)}\sim f_{+}$ where
$\left(\frac{1}{\Phi_{0}^{2}(m_{\phi}^{2}-\omega_{0}^{2})^{2}}\frac{\text{d}U}{\text{d}f}-\left(\Omega-\alpha A\right)^{2}f\right)_{{f=f_{+}}}=0.$ (13)
For the specific potential (9), we have $f_{+}\sim 1$.
In this region, the gauge field is massive and satisfies
$\nabla^{2}A=\alpha f_{+}^{2}(A\alpha-\Omega)$ (14)
We will take $f_{+}$ to be approximately constant. In that case, the solution for the gauge field is
$\alpha A^{(0)}=\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}$ (15)
In the exterior $\rho\gg R$, the scalar is extremely close to zero. In this region, the solution for the gauge field is
$\alpha A^{(0)}={Q_{out}\over\rho}$ (16)
Continuity of the field and its derivative at $\rho=R$ then implies
@@
The transition region $\rho\sim R$ is the most interesting. In this region, the scalar varies quickly from $f_{+}$ to zero over a range of order 1. The gauge field on the other hand varies much more slowly (for small $\alpha$) and can be taken to be the matching value of the solution (15, 16, 17). We therefore take the gauge field to be approximately constant with the value
@@
The equation for $f$ is now
$(f^{(0)})^{\prime\prime}+\frac{2}{R}(f^{(0)})^{\prime}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-(\Omega_{R}^{(0)})^{2}\right)$ (19)
where we have replaced the explicit factor of $\rho$ by $R$, and
@@
Since $\Omega_{R}^{(0)}$ is being treated as effectively constant, this maps the problem to a global Q-ball equation where $\Omega_{R}^{(0)}$ plays the role of the frequency. Since the frequency and the radius are related for a global Q-ball, this equation can only be solved if $\Omega_{R}^{(0)}=\Omega_{e=0}$, where $\Omega_{e=0}$ is the value of $\Omega$ which would produce a global Q-ball of radius $R$.
Following Heeck:2021zvk , the solution of the equation (19) is
@@
with
@@
This can be shown to be an accurate description of the Q-ball profile Heeck:2021zvk .
@@ IV Perturbed Q-balls
IV.1 Ansatz for the perturbation
We now consider a perturbation of the Q-ball where the surface becomes nonspherical; we take the surface to be at
@@ IV.1 Ansatz for the perturbation
We will below, for conciseness, denote the Legendre polynomials $P_{L}(\cos(\theta))$ as $P_{L}$.
We modify the transition profile to reflect the perturbation; we can write the ansatz for the new profile as
$f(\rho)=\frac{f_{+}}{\sqrt{1+2e^{2(\rho-R(\theta))}}}\,.$ (24)
Typically, such a perturbation will not lead to a solution of the equations of motion, and the perturbation will either grow or decay. In either case, $\epsilon$ will have a time dependence. Even if we have a perturbation which is initially static (i.e. the initial time derivative $\dot{\epsilon}=0$), the dynamics will drive the value of $\epsilon$ away from its initial value. That is, the dynamics induces a nonzero value of $\ddot{\epsilon}$. Instead of a constant $\epsilon$, therefore, we can set (for small times) $\epsilon(t)=\epsilon(1+{1\over 2}\gamma t^{2})$. If $\gamma<0$, the perturbation is presumably stable, while if $\gamma>0$, the perturbation grows, and therefore is unstable. We will assume however that $\gamma$ is small, so that we can approximately treat $\epsilon$ as constant.
We now look for an equation that determines $\gamma$.
IV.2 The gauge field
The first step is to calculate the profile of the gauge field, and in particular the value of the gauge field in the transition region.
Following the previous section, we use the thin wall approximation to find the gauge field. In the interior we will again have $f=f_{+}$ while in the exterior $f=0$. The equations of motion for the gauge field in the interior and exterior are therefore unchanged.
At the linearized level, the gauge field will obtain harmonics proportional to $P_{L}$. The solution will be therefore of the form
$\alpha A=\left\{\begin{array}[]{c}\Omega+Q_{in}{\sinh(\alpha f\rho)\over\rho}+\epsilon P_{in}F_{L}(\alpha f\rho)P_{L}\qquad\rho\ll R(\theta)\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}Q_{out}{1\over\rho}+\epsilon P_{out}{1\over\rho^{L+1}}P_{L}\qquad~{}~{}~{}~{}~{}~{}~{}~{}\rho\gg R(\theta)\end{array}\right.$ (27)
where $F_{L}(\alpha f\rho)=j_{L}(i\alpha f\rho)$, and $j_{L}$ is a spherical Bessel function. We list the first few $F_{L}$ in an appendix.
We again require the gauge field and its derivative to be continuous at the boundary $\rho=R(\theta)$. This yields
@@ IV.2 The gauge field
This then implies that at $\rho=R(\theta)$, we have
$\alpha A_{R}=\alpha A^{(0)}_{R}+\epsilon P_{L}a_{R}$ (29)
where
$a_{R}={\Omega F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\Omega\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)$ (30)
IV.3 The scalar equation
Once again we note that the gauge field is slowly varying as a function of $\rho$, and can be taken to have a value $A_{R}(\theta)$ over the transition region. We then have
$-\ddot{f}+f^{\prime\prime}+\frac{2}{\rho}f^{\prime}+{1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (31)
where $\Omega_{R}=\Omega-\alpha A_{R}(\theta)$.
In the ansatz (24), the time dependence and the angular dependence are linked to the radial dependence. Indeed, we can write
$-\ddot{f_{T}}={f_{T}}^{\prime}R\gamma{\epsilon}P_{L}$ (32)
and similarly
${1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f_{T})=f_{T}^{\prime}(\epsilon){1\over R}L(L+1)P_{L}$ (33)
Note that this is small, and so to leading order we can treat the scalar field as only depending on $\rho$.
Substituting these into equation (31), we find that in the transition region, the equation is
$f^{\prime\prime}+\frac{2}{R_{eff}}f^{\prime}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (34)
with
$\frac{2}{R_{eff}}=\frac{2}{R(\theta)}+R\gamma\epsilon P_{L}+\epsilon{1\over R}L(L+1)P_{L}$ (35)
The equation (34) is identical to the equation (19)in the previous section with the replacement $\Omega^{(0)}_{R}\to\Omega_{R},R^{(0)}\to R_{eff}$ The relation (22) is therefore modified at large radius to
@@ IV.3 The scalar equation
Eliminating $R_{eff}$, we find
$2(\Omega_{R}^{2}-{\Omega_{0}^{2}})=2((\Omega^{(0)}_{R})^{2}-{\Omega_{0}^{2}})+R\gamma\epsilon P_{L}+\epsilon{1\over R}(L(L+1)-2)P_{L}$ (37)
This implies
$-4\Omega^{(0)}_{R}a_{R}-{(L+2)(L-1)\over R}=R\gamma$ (38)
where $a_{R}$ is defined in (29).
IV.4 The prediction for the instability
Combining (30) and (38), we have
$R^{0}\gamma=-4\Omega\Omega^{(0)}_{R}\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)\right)-{(L+2)(L-1)\over R}$ (39)
where
$\Omega=\Omega_{R}^{(0)}{\alpha f_{+}R\over{\tanh(\alpha f_{+}R)}}\qquad\Omega^{(0)}_{R}=\left(\sqrt{\Omega_{0}^{2}+{1\over R^{2}}}\right)$ (40)
This is our final result for $\gamma$ in terms of the other parameters. As we have said above, $\gamma>0$ indicates an instability.
@@ V Comparison with Numerics
We now compare to the numerical results of  Kinach:2022jdx .
V.1 The point P2
The authors of Kinach:2022jdx consider the sextic potential with
$V=|\phi|^{2}-(1/2)|\phi|^{4}+{0.2\over 3}|\phi|^{6}$ (41)
They analyzed in detail a specific parameter choice point (which they denote as P2), with $\omega=0.9958,e=0.02,r=65$ in their conventions. After rescaling to match our conventions, we find that this parameter point corresponds to
$\Omega_{0}^{2}=1/15\qquad\Omega=1.028\qquad R=62.9\qquad\alpha=0.057\qquad\alpha R=3.56$ (42)
From (40), we also have
$\Omega^{(0)}_{R}=\left(\sqrt{1/15+{1\over 62.9^{2}}}\right)=0.2587$ (43)
We now calculate the value of $R\gamma$ for each $L$. The results are shown in Table 1.
L 1 2 3 4 5 6
$R\gamma$ 0 0.123 0.148 0.102 -0.0001 -0.15
Table 1: $R\gamma$ as a function of $L$ for the point P2
We first note that $L=1$ appears to be a mode which is marginal i.e. neither stable nor unstable. This mode in fact corresponds to a translation of the Q-ball, and will be henceforth ignored.
We see from the above that $L=2,3,4$ all have $\gamma>0$, and hence are all predicted to be unstable modes. $L=5$ is apparently stable, but the value is very close to zero, and this result may be changed in a more accurate analysis. Other modes are stable.
This is to be compared with the results of Kinach:2022jdx , which finds that the Q-ball with this parameter choice is unstable, with an instability mode with $L=4$. They also find the instability is near the surface of the Q-ball. This is all consistent with our results. The only discrepancy is that we predict an $L=3$ mode which is in some sense more unstable; however, the absence of this mode, while somewhat puzzling, may plausibly be attributed to some feature of the numerical analysis.
V.2 Location of the Instability Threshold
We now look at the predicted instability threshold for $\alpha=0.057$, i.e. the first value of $R$ for which some $\gamma$ becomes positive. To do this, we calculate $R\gamma$ as a function of $R$. For $\Omega^{(0)}_{R}$, we use
@@ V.2 Location of the Instability Threshold
For $\Omega$, we scale this from the point P2 as
${\Omega\over 1.028}={\Omega_{R}^{(0)}\over 0.2587}{\alpha f_{+}R\over{\tanh(\alpha f_{+}R)}}{\tanh(3.26)\over 3.26}$ (45)
We then find that the smallest radius of instability occurs for $L=2$, and for a predicted value of $\alpha R=1.99$. We therefore predict that all radii above this value are unstable.
We compare this to the radius where ${d\Omega\over dR}$ becomes positive. This is at $\alpha R=0.51$, and it can then be checked straightforwardly, that for the radii which are in the unstable regime, we indeed have ${d\Omega\over dR}>0$.
However, it is to be noted that here we predict a significant difference between the instability threshold and the location where ${d\Omega\over dR}>0$. This is not in agreement with the results of Kinach:2022jdx , where these two appear to be at very similar values. We discuss this further in the next section.
V.3 Larger couplings: the point P1
The second parameter point discussed in Kinach:2022jdx is denoted P1, and has a much larger value of $e=0.17$, corresponding to a value of $\alpha=0.481$.
For such large values of $\alpha$, the analysis of the previous section cannot be trusted. Nevertheless, we can see what our formula predicts. Following the same procedure as the previous subsection, we find that the instability threshold is predicted to occur for $L=2$ at a value $\alpha R=6.7$. This corresponds to a predicted threshold radius of about 14. The Q-ball radius is not given in Kinach:2022jdx , but general considerations Loginov:2020xoj show that $R$ should scale as $\alpha^{-1}$ and should therefore be about 8. This indicates that the instability has not yet been reached, and P1 is stable. This is in agreement with the numerical results.
V.4 Smaller couplings
On the other hand, we can ask what happens if the coupling is taken to be smaller. We find that the value of $\alpha R$ at the threshold decreases. For example if we take $\alpha=10^{-4}$, we find a threshold radius (i.e. the radius where $\gamma=0$) that satisfies $\alpha R=0.21$. Numerical analysis of these very small $\alpha$ will shed light on whether our formulae are accurate for all Q-balls.

### v3

@@ I Introduction
The stability of these solitons has been discussed in many papers, and the picture is still unclear. For global Q-balls, it has been shown Lee:1991ax ; Friedberg:1976me ; Paccetti:2001uh that the Q-balls are stable as long as ${d\omega\over dQ}<0$ (these quantities are defined below). For gauged Q-balls, the analogous statement is known to not be true Panin:2016ooo .
Recently, a numerical analysis of the stability of gauged Q-balls was carried out Kinach:2022jdx . It was found that gauged Q-balls were unstable for certain parameters, and in particular were unstable to modes which had an angular dependence. On the other hand, for other choices of parameters, the corresponding Q-balls were stable. In general, Q-balls with ${d\omega\over dQ}<0$ were stable.
@@
We then show in section V that the instability that we predict agrees in many details in common with the numerical analysis of Kinach:2022jdx . In particular, we indeed find that for one parameter choice in Kinach:2022jdx , the $L=4$ mode is unstable, in agreement with the numerical results. Similarly, we find that for another parameter choice, the instability is absent, in agreement with the numerical analysis.
@@ II Q-Balls: Overview and Notation
We review here the basic properties of Q-balls; we will follow the notation of Heeck:2020bau ; Heeck:2021zvk .
@@ III Spherical gauged Q-balls
We begin by reviewing the construction of spherical Q-balls. This section follows the analysis of Heeck:2021zvk .
The spherical Q-ball is approximately a sphere of radius $R$ with a constant internal density. More precisely, in the interior, the scalar field is approximately at the false vacuum solution $f=f_{+}$ (where $f_{+}$ is the nonzero solution of the equation 5) and at the Q-ball surface, the field rapidly transitions to the true vacuum $f=0$. The scalar profile is thus to a first approximation
$f^{(0)}=\left\{\begin{array}[]{c}f_{+}\qquad\rho<R\\ 0\qquad\rho>R\end{array}\right.$ (12)
The gauge field is a nondynamical field, and is determined from the scalar charge distribution through the Maxwell equation 6. For the scalar profile above, the solution for the gauge field is
$\alpha A^{(0)}=\left\{\begin{array}[]{cc}\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho<R\\ {Q_{out}\over\rho}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\rho>R\end{array}\right.$ (15)
Note that the typical scale of variation of $A$ is inversely proportional to $\alpha$.
Near $\rho\sim R$, the scalar field is varying very quickly over a short distance of order 1, which we call the transition region. However, since the gauge field varies much more slowly than the scalar, we can take the gauge field to be approximately smooth over the transition region. Continuity of the gauge field and its derivative at $\rho=R$ then implies
@@
We note that the value of the gauge field at $\rho=R$ is
@@
We can now require consistency between the obtained gauge field and the initial ansatz for the scalar profile. To do this, we solve the equation for the scalar field in the gauge field background (15). This equation is now
${d^{2}f^{(0)}\over d\rho^{2}}+\frac{2}{\rho}{df^{(0)}\over d\rho}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-\left(\Omega-\alpha A^{(0)}\right)^{2}\right)$ (18)
This is particularly important to solve in the transition region, where the scalar field varies quickly from the false vacuum to the true vacuum. This requires a very precise relation between the potential parameters and the radius of the Q-ball, as discussed in Heeck:2021zvk for the case of global Q-balls, and which we now review.
In the gauged Q-balls case that we are discussing here, the scalar field in the transition region would satisfy the equation
${d^{2}f^{(0)}\over d\rho^{2}}+\frac{2}{R}{df^{(0)}\over d\rho}=f^{(0)}\left(1-4(f^{(0)})^{2}+3(f^{(0)})^{4}+{\Omega_{0}^{2}}-(\Omega_{R}^{(0)})^{2}\right)$ (19)
where we have replaced the explicit factor of $\rho$ by $R$, and set the gauge field to be its value at $\rho=R$. This has allowed us to replace $\left(\Omega-\alpha A^{(0)}\right)$ by the constant
@@
This maps the problem to the form of a global Q-ball equation, and we can use the methods in  Heeck:2021zvk to solve it.
We can treat the term $\frac{2}{R}{df^{(0)}\over d\rho}$ as a small frictional term. In its absence, there is an exactly conserved energy, and this allows the equation to be solved exactly; consistency of energy conservation implies that this is a solution in the limiting case when the false vacuum and the true vacuum have the same energy. In this limit, the solution of the equation (19) is
@@
If the false vacuum and the true vacuum have slightly different energies, the frictional term must be included if the solution is to exist, and furthermore $R$ must have a precise value. Specifically, the change in the energy between the two vacua should be balanced by the effect of the frictional term. The details of the derivation of the resulting consistency condition are in Heeck:2021zvk and show that the radius must satisfy
@@
This can be shown Heeck:2021zvk to be an accurate prediction for the Q-ball radius.
@@ IV Perturbed Q-balls
We now look for potential instabilities of these Q-balls.
We do these by considering deformations of the spherical Q-ball. The deformation of the Q-ball will in general not lead to a solution of the equations of motion. The deformations may grow, indicating that the Q-ball is unstable, or they may decay away. However, if there is a transition between stable Q-balls and unstable Q-balls, then exactly at threshold, the deformation will neither grow nor decay, and therefore the deformed Q-ball will be a solution of the equations of motion.
Now the spherical Q-ball has an approximately constant density in the interior, and zero density outside. We will consider perturbations of the Q-ball where the density is assumed to be approximately the same constant in the interior, but where the surface deforms in some way. It is likely that other deformations (e.g. where the internal density is nonconstant) are possible; we leave the search for these to future work. The assumptions we make will hence lead to a conservative estimate for the region of stability (i.e. our analysis will not find the most general instability, but any Q-balls that we find to be unstable will indeed be unstable).
It is fairly immediate that under these assumptions on the perturbations, a spherical perturbation would change the charge. We must therefore consider perturbations with an angular dependence, such that the total charge does not change to leading order.
At this point, then, we are led to consider a perturbation where the Q-ball surface is deformed to be at
@@
$P_{L}(\cos(\theta))$ is a Legendre polynomial; we shall denote this as $P_{L}$, suppressing the $\theta$ dependence. $\epsilon$ parametrizes the size of the perturbation, and can be taken to be parametrically small.
We now follow the same approach as the previous section to construct these Q-balls.
First, we will make an initial ansatz that the scalar field is at the false vacuum in the interior and zero outside i.e.
$f=\left\{\begin{array}[]{c}f_{+}\qquad\rho<R(1+\epsilon P_{L})\\ 0\qquad\rho>R(1+\epsilon P_{L})\end{array}\right.$ (26)
The equation of motion 6 for the gauge field can then be solved to find
$\alpha A=\left\{\begin{array}[]{c}\Omega+Q_{in}{\sinh(\alpha f_{+}\rho)\over\rho}+\epsilon P_{in}F_{L}(\alpha f_{+}\rho)P_{L}\qquad\rho<R(1+\epsilon P_{L})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}Q_{out}{1\over\rho}+\epsilon P_{out}{1\over\rho^{L+1}}P_{L}\qquad~{}~{}~{}~{}~{}~{}~{}~{}\rho>R(1+\epsilon P_{L})\end{array}\right.$ (29)
where $F_{L}(\alpha f_{+}\rho)=j_{L}(i\alpha f_{+}\rho)$, and $j_{L}$ is a spherical Bessel function. We list the first few $F_{L}$ in an appendix.
We again require the gauge field and its derivative to be continuous at the boundary. This yields
@@
At the boundary, the value of the gauge field is now
$\alpha A_{R}=\Omega\left(1-{1\over\alpha f_{+}R}{\tanh(\alpha f_{+}R)}\right)+\epsilon P_{L}\Omega\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-1+{\tanh(\alpha fR)\over\alpha fR}\right)$ (32)
We can now require again consistency between the obtained gauge field and the initial ansatz for the scalar profile. To do this, we solve the equation for the scalar field in the gauge field background (29). We focus on the transition region, where the scalar field varies quickly from the false vacuum to the true vacuum. Once again we note that the gauge field is slowly varying as a function of $\rho$, and can be taken to have a value $A_{R}$ over the transition region.
The differential equation for the scalar is
${\partial^{2}f\over\partial\rho^{2}}+\frac{2}{\rho}{\partial f\over\partial\rho}+{1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (33)
where $\Omega_{R}=\Omega-\alpha A_{R}$. Since $f$ is now dependent on $\theta$, we have kept the angular derivatives in the Laplacian.
This is now a differential equation in two variables, which is obviously much more complex. However, we note that all the radial and angular dependence in the ansatz (26) is at the surface of the Q-ball, and hence these dependences are related. To make this precise, note that the ansatz (26) only depends on $\rho-R\epsilon P_{L}$. We therefore have
${1\over\rho^{2}\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta}f)={df\over d\rho}{1\over R}L(L+1)\epsilon P_{L}$ (34)
where we have only kept the terms linear in $\epsilon$. The angular derivative terms therefore yield terms proportional to ${df\over d\rho}$.
In the transition region, we can also set $\rho=R(\theta)=R(1+\epsilon P_{L}(\cos(\theta)))$. The scalar equation then becomes
${\partial^{2}f\over\partial\rho^{2}}+\left(\frac{2}{R(\theta)}+{1\over R}L(L+1)\epsilon P_{L}\right){\partial f\over\partial\rho}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (35)
The partial differential equation now only has derivatives with respect to $\rho$. This can then be solved separately for each $\theta$. For example, one can fix $\theta$, and then find the profile by starting at $f=f_{+}$ at $\rho=0$ and stepping outward to large $\rho$. We therefore get a profile for each value of $\theta$.
For each value of $\theta$, the equation is again of the form of a global Q-ball equation. We can write the equation (35) as
${d^{2}f\over d\rho^{2}}+\frac{2}{R_{eff}}{df\over d\rho}=f\left(1-4f^{2}+3f^{4}+{\Omega_{0}^{2}}-\Omega_{R}^{2}\right)$ (36)
where we have defined
$\frac{2}{R_{eff}}=\frac{2}{R(\theta)}+\epsilon{1\over R}L(L+1)P_{L}$ (37)
This is identical to the equation (19)in the previous section with the replacement $\Omega^{(0)}_{R}\to\Omega_{R},R\to R_{eff}$. The solution may therefore be read off from (22) and states that the perturbed Q-ball satisfies
@@
This may be simplified by matching terms linear in $\epsilon$ on both sides, and yields the equation for a stable deformed Q-ball as
${\Omega F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\Omega\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)=-{(L+2)(L-1)\over 4\Omega^{(0)}_{R}R}$ (39)
This is also the relation that gives the instability threshold. That is, if there is a transition from a region of stable Q-balls to a region of unstable Q-balls, the boundary is set by this relation.
What if this relation is not satisfied? Then the deformed Q-ball does not satisfy the equations of motion, and it will either relax back to the spherical Q-ball, or the deformation grows, indicating an instability. At least close to the threshold, we can track this by giving $\epsilon$ a small time dependence. Following the same derivation, we find
${R\ddot{\epsilon}\over\epsilon}=-4\Omega\Omega^{(0)}_{R}\left({F_{L}(\alpha fR)\tanh(\alpha fR)\over F^{\prime}_{L}(\alpha f\rho)+F_{L}(\alpha fR){(L+1)\over\alpha fR}}-\left(1-{\tanh(\alpha fR)\over\alpha fR}\right)\right)-{(L+2)(L-1)\over R}$ (40)
Therefore, if the right-hand side of the above equation is negative, the Q-ball is stable, but if it is positive, there is a instability.
@@ V Comparison with Numerics
We will now compare this theoretical prediction for the instability threshold of gauged Q-balls to the numerical results of  Kinach:2022jdx .
The authors of Kinach:2022jdx consider Q-ball instabilities in several types of potentials; one of the models they consider is a gauged Q-ball in a sextic potential with
$V=|\phi|^{2}-{1\over 2}|\phi|^{4}+{0.2\over 3}|\phi|^{6}$ (41)
For this potential they present three results:
(i) For a particular parameter point P2 (details below) the Q-ball is unstable. The instability has $L=4$.
(ii) For a particular parameter point P1 (details below) the Q-ball is stable.
(iii) For a particular choice of parameters, they find the instability threshold.
We review these results, and compare to our prediction.
(i) The parameter point P2 has $\omega=0.9958,e=0.02,r=65$. After rescaling to match our conventions, we find that this parameter point corresponds to
$\Omega_{0}^{2}=1/15\qquad\Omega=1.028\qquad R=62.9\qquad\alpha=0.057\qquad\Omega^{(0)}_{R}=0.2587$ (42)
For these parameters, we calculate the right-hand-side of eqn (40) for each value of $L$. We find that the right-hand-side vanishes for $L=1$; this reflects the translational mode of the Q-ball, and must be ignored. The right-hand-side is positive for $L=2,3,4$ and hence all these modes are predicted to lead to an instability. The value of the RHS for $L=5$ is negative, but very close to zero, and this result may be changed in a more accurate analysis. For $L>5$, the right-hand-side is negative, and hence these modes do not lead to an instability .
This is to be compared with the results of Kinach:2022jdx , which finds that the Q-ball with this parameter choice is unstable, with an instability mode with $L=4$. They also find the instability is near the surface of the Q-ball. This is all consistent with our results.
(ii) The parameter point P1 discussed in Kinach:2022jdx has $e=0.17,\omega=0.9976$. This corresponds in our notation to $\Omega=1.029,\alpha=0.481$. The value of $\Omega_{0}$ is again $1/\sqrt{15}$.
For such large values of $\alpha$, the analytical prediction of the previous section cannot be trusted. Nevertheless, we can see what the formula predicts. We find that the lowest instability threshold (i.e. the smallest unstable $R$) is predicted to occur for $L=2$ at a value $\alpha R=6.7$. This corresponds to a predicted threshold radius of about 14 (that is, Q-balls with a radius smaller than 14 are predicted to be stable, while larger Q-balls are predicted to be unstable).
The Q-ball radius at parameter point P1 is not given in Kinach:2022jdx , but we may compare the charges at these two parameter points (the charge for P1 is given as 387.5 while the charge for P2 is given as $1.480\times 10^{6}$) to estimate the ratio of the raidii for these two parameter points. This shows that the radius at parameter point P1 should be about 4. Our prediction would therefore be that the Q-ball at parameter point P1 is stable. This is in agreement with the numerical results.
(iii) For $e=0.02$ (corresponding to $\alpha=0.057$ in our conventions), the authors of Kinach:2022jdx also present the value of the instability threshold i.e. the first value of $R$ for which the Q-ball is unstable. This is found to be very close to the radius where ${d\Omega\over dR}$ becomes positive. This is at $\alpha R=0.51$.
Let us see what the equation (40) predicts. Unfortunately, we do not have the other parameter values of the numerical calculation, and we are forced to make estimations of these. Specifically, we need to find $\Omega,\Omega^{(0)}_{R}$ at the threshold.
We first calculate $\Omega^{(0)}_{R}$ using (22) to find
@@
For $\Omega$, we use the known values at parameter point P2, as well as equation (20) to find $\Omega$ as
${\Omega\over 1.028}={\Omega_{R}^{(0)}\over 0.2587}{\alpha f_{+}R_{t}\over{\tanh(\alpha f_{+}R_{t})}}{\tanh(3.26)\over 3.26}$ (44)
Substituting $\Omega,\Omega^{(0)}_{R}$ into equation (40) gives us an equation determining the instability threshold for each $L$. We find that the smallest radius of instability occurs for $L=2$, and for a predicted value of $\alpha R=1.99$. We therefore predict that all radii above this value are unstable.
Here we have a significant difference between our predicted instability threshold of $\alpha R=1.99$, and the numerically calculated threshold Kinach:2022jdx of $\alpha R=0.51$. We discuss this further in the next section.